OpenStudy (anonymous):

find the derivative of this function \[y=xe ^{-6x}\]

OpenStudy (anonymous):

using integration by parts, let u = x and dv = e^(-6x)

OpenStudy (anonymous):

o whoops derivative

OpenStudy (anonymous):

yes but how do i derivative the e^-6x

OpenStudy (anonymous):

use the product rule

OpenStudy (anonymous):

e^(-6x) + x(-6)e^(-6x)

OpenStudy (anonymous):

since the product rule says d/dx f(x)g(x) = f'(x)g(x) + g'(x)f(x) so let f(x) = x and g(x) = e^(-6x)

OpenStudy (anonymous):

so thats the answer? e^(-6x)?

OpenStudy (anonymous):

have look at the "solution.gif" above.

OpenStudy (anonymous):

i did but i just dont understand....

OpenStudy (anonymous):

the derivative is right hand side of the equation.

OpenStudy (anonymous):

"d/dx" means the derivative with respect to variable x.

OpenStudy (anonymous):

i know but when i submit into my web assign it said its wrong...i dont understand how i did it myself too

OpenStudy (anonymous):

i think result should be (e^(-6*x))*(1-6*x) * means multiplication.

OpenStudy (anonymous):

ok..let me see